APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday December 2, 2003 (MJD 52976) ---=== OBSERVING TEAM ===--- Atsuko Nitta (swing) Dan Long Mike Harvanek Scot Kleinman Connie Rockosi Jim Gunn ---=== OBSERVING PLAN ===--- Shake and science. ---=== OBSERVING SUMMARY ===--- Shake: ------ Changes in procedure were verified. tccMon now has a combined count up timer/total integration time field and the scale factor field has been restored. The SOPGUI exposure countdown timer now sounds a 4 beep alarm 5 minutes before the end of each exposure. Imager flats were taken and a procedure was established for obtaining them on a monthly basis. Progress was made on the new guider code but it is not ready for prime time and will require more shaking. Poor seeing prevented us from obtaining donuts for JEG's imager tests. Science: -------- Finished 2 plates (1301, 1591). Another plate (1665) may also be finished. Although SoS says otherwise, we are suspicous of the reductions. Could someone please have a look at the data for this plate (mjd 52975, 52976) and tell us if we can declare it done. Thanks. Also, 2 days ago we asked for someone to look at the data for plate 1253 (mjd 52974) to see if we could declare it done and we still have not heard anything back. This is the end of the dark run and it would be nice to know if this plate is done so the plate can be unplugged and the fibers relaxed. Could someone PLEASE have a look and get back to us ASAP. ---=== OBSERVING LOG ===--- Afternoon --------- Due to on-going work Connie was doing, I did not reboot host this afternoon. Spectros did not need any adjustments. Everything else checked out fine. Night ----- We began with shake up. We have new versions of IOP/SOP, tccmon and procedures. All changes in the procedures changeLog were verified. The new version of tccmon has combined the count up timer and the total integration time into a single field called GCTimer. The new value is expressed as a fraction with the count up timer as the numerator and the total integration time as the denominator. Also, the scale factor field has returned in this new tccmon. In the new version of SOP, the SOPGUI exposure countdown timer sounds an alarm 5 minutes before the end of each exposure. The alarm is 4 beeps and the sound of the beep is different from that of the standard beep of each computer. Imager flats were taken and a procedure was established for obtaining them. Imager flats should be taken once a month during shake up. The enclosure should be as dark as possible (e.g., push in a stop button in the lower enclosure and release all those in the upper enclosure). The drift should be started approximately 30 minutes after sunset (not earlier) and should run for about 30 frames. The new guider code was then tested. It was found to guide properly and it did not crash SOP. However, some of the bells and whistles were still inoperative (bogus efficiency meter, failure of the initial guiderFocusLoop). While Connie continued to work on the new guider code, we proceeded to do science (spectroscopy) while continuing to test this new code. Plate 1665 (cartridge 2): Guide stars on the gotoField. This was actually shake time and we were testing the new guider code while taking observations. Consequently, the time tracking is not very accurate. There are large periods of what appear to be inactivity and inordinately large setup times that are due to the testing. We also lost a 20 minute science exposure when a hybrid version of SOP (new guider code accidentally joined with the broken shake version of sop v3_122_0) crashed. The first 3 exposures were all good. Exposures 4, 5 and 6 had s/n^2 < 2 for b2 and r2 but b1 and r1 were good. The 7th exposure had s/n^2 < 2 for r2 only and the rest were good. There was something suspicious about the poor signal in camera 2 for the 6th and 7th exposures, however. Between exposures 5 and 6, the signal in camera 1 nearly doubled while there was no corresponding increase in camera 2. Ditto for exposure 7. Perhaps a whopping fiber (or something else) is causing SoS to have problems with the reductions. Could someone please have a look at the data for this plate and tell us if we can declare it done. Thanks. The moon set but thin clouds prevented imaging. We could not even take donuts for JEG's imager tests because of poor seeing so we continued with spectroscopy. Plate 1301 (cartridge 5): An FK5 star was needed to acquire the field. However, more problems with the new guider code (reinitialization problem associated with loading a new cartridge and fitting for dark subtraction as a function of exposure time) delayed science. When combined with previous data (52974-75), this plate was finished in a single 1500 s exposure. Plate 1591 (cartridge 1): Guide stars on the gotoField. This plate passed within 1.3 degrees of zenith. One of the science exposures (23729) was paused for about 8-10 minutes because of the trouble guiding this close to zenith. The watcher notified us of max acceleration exceeded errors in both azimuth and rotator during this plate. The plate was finished in 4 exposures (4500 s total). endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 4300 22:38Z 22:58Z 100 O 120.21 125.25 48 ignore 4301 23:57Z 00:08Z 100 O 139.99 142.78 33 bias on telescope 4302 00:10Z 01:27Z 100 O 143.16 162.49 143 ignore 4303 01:30Z 01:50Z 100 O 163.24 168.25 46 calibration flats ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): Yes QA Procedures Done (y/n): Yes UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 23696, plate -9999 ------- 17:15 23695 0.0 bias 22:33 23696 0.0 bias ----- sequence 23697, plate 1301 ------- 22:41 23697 10.0 flat test 22:44 23698 2.0 arc test 22:47 23699 2.1 arc test 22:49 23700 2.1 arc test ----- sequence 23701, plate -9999 ------- 00:05 23701 0.0 bias 00:12 23702 0.0 bias 00:40 23703 0.0 bias 00:47 23704 0.0 bias 00:51 23705 0.0 bias ----- sequence 23706, plate 1301 ------- 01:59 23706 65.7 target testing - declared bad ----- sequence 23707, plate 1665 ------- 04:07 23707 10.0 flat 28.6 25.4 18.0 8.8 tonight 04:09 23708 2.0 arc 17.3 15.0 21.6 10.8 52975 04:33 23709 1200.1 target ------------------- 04:54 23710 1100.1 target 45.9 40.4 39.6 19.6 ----- sequence 23711, plate -9999 ------- 05:01 23711 0.0 bias 05:04 23712 0.0 bias ----- sequence 23713, plate 1665 ------- 05:34 23713 1100.1 target ----- sequence 23714, plate -9999 ------- 06:24 23714 0.0 bias ----- sequence 23715, plate 1665 ------- 07:01 23715 1200.1 target b2, r2 < 2 07:24 23716 1200.1 target b2, r2 < 2 ----- sequence 23717, plate -9999 ------- 07:36 23717 0.0 bias ----- sequence 23718, plate 1665 ------- 07:51 23718 0.0 bias 08:29 23719 1200.2 target b2, r2 < 2 08:51 23720 1200.1 target r2 < 2 08:55 23721 10.0 flat 08:57 23722 2.0 arc ----- sequence 23723, plate 1301 ------- 5.4 4.2 6.7 3.9 tonight 09:55 23723 10.0 flat 16.6 8.1 18.3 7.7 52975 09:57 23724 2.0 arc 5.0 4.4 5.4 4.2 52974 10:25 23725 1500.1 target ------------------- 27.0 16.7 30.4 15.8 DONE ----- sequence 23726, plate 1591 ------- 10:48 23726 10.0 flat 10:51 23727 2.0 arc 11:09 23728 900.1 target 26.2 16.4 29.5 16.2 DONE 11:45 23729 1200.1 target 12:08 23730 1200.1 target 12:31 23731 1200.1 target 12:34 23732 10.0 flat 12:36 23733 2.0 arc ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 04:06Z 2 1665 122.41 -0.0081 69.93 0.0010 -71.48 0.0103 1.000290 07:21Z 2 1665 239.56 -0.0102 67.36 0.0045 75.64 0.0249 1.000240 09:54Z 5 1301 48.43 -0.0029 57.08 -0.0006 219.45 -0.0138 1.000240 10:47Z 1 1591 83.64 -0.0112 82.05 -0.0005 258.68 0.0269 1.000220 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL7074 Stays: JL7075 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 04:05Z 2 1665 1.00029 -2863 -989 250 122 69.9 5.3 17 29 ---- 1.8" 07:22Z 2 1665 1.00024 -2369 -684 142 240 67.3 3.0 11 76 ---- 2.4" 09:53Z 5 1301 1.00024 -2369 -721 80 49 56.9 3.4 8 150 ---- 2.6" 10:47Z 1 1591 1.00022 -2171 -601 80 84 81.9 3.1 8 178 ---- 1.9" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:09Z 47 8 12 312 (NW) 249 - 22:41Z 47 8 17 318 (NW) 203 - 23:13Z 46 11 4 293 (WNW) 310 - 23:45Z 45 10 12 346 (NNW) 281 - 00:13Z 44 11 9 357 (NNW) 327 - clear 01:04Z 43 12 23 334 (NNW) 677 - " 01:35Z 43 6 11 10 (N) 303 - " 02:06Z 42 8 9 343 (NNW) 388 - " 02:38Z 41 8 17 6 (N) 480 - " 03:11Z 40 6 25 5 (N) 429 - " 03:42Z 41 -1 21 12 (NNE) 318 - " 04:13Z 41 -3 19 24 (NNE) 259 - light clouds 04:43Z 41 -1 14 13 (NNE) 302 - " " 05:13Z 40 -3 17 3 (N) 279 - " " 05:43Z 39 -1 18 14 (NNE) 208 - " " 06:13Z 38 -1 26 30 (NNE) 201 - " " 06:44Z 38 -3 25 33 (NNE) 127 - " " 07:14Z 38 -7 14 67 (ENE) 131 - " " 07:45Z 43 -4 6 88 (E) 58 - " " 08:16Z 43 -4 1 179 (S) 28 - " " 08:46Z 40 -6 8 179 (S) 49 - " " 09:17Z 39 -7 14 149 (SSE) 173 - clear 09:47Z 38 -7 11 141 (SE) 204 - light clouds 10:17Z 37 -8 9 161 (SSE) 152 - " " 10:48Z 37 -9 8 167 (SSE) 146 - " " 11:18Z 39 -6 7 180 (S) 226 - " " 11:50Z 38 -7 3 117 (ESE) 159 - " " 12:20Z 35 1 6 78 (ENE) 266 - " " 12:51Z 35 1 5 92 (E) 317 - " " ---=== TELESCOPE STATUS ===--- 23:15Z Enclosure doors opened, fans and blowers on, louvers opened. 02:40Z Enclosure off (delayed by shakeup). 12:50Z Enclosure on, doors closed, fans and blowers off, louvers closed. Telescope stowed at 30 degrees with cartridge 9 (plate 1180) mounted. Counterweights at 280. Spectro auto-fill connected. Imager dewar 13:05 162 lb Spectro dewar 13:05 261 lb No interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v3_123_0 Watcher: v2_24_0 MCP: v5_23_0 TPM: tpm_v2_36_0 AstroDa: v14_47 TCC: 2.7.1 sdssProcedures: v1_80 tccMon: v2_6 SoS: v4_10_6 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.5500 0.5450 Axial B 0.6310 0.6370 Axial C 1.0880 1.0850 Trans D -8.4600 -8.4550 Lateral E 10.7950 10.7442 Lateral F 10.7950 10.7823 GALILS Commanded: 4600. -2850. 850. -9250. 18400. 17850. Actual: 4600. -2866. 866. -9275. 18411. 17825. SETMIR VALUES PriDesOrient: 0.00 0.00 15.20 829.90 1.60 PriOrient: 0.00 0.00 14.91 830.49 3.19 SECONDARY: ---------- Focus: 0.00 Air Temp.: 8.3 Alt.: 30.206303 MIGS TONIGHT NOMINAL Axial A 1.5160 1.5150 Axial B 1.1090 1.1100 Axial C 1.1160 1.1150 Trans D -0.8260 -0.8260 GALILS Commanded: 1628814. 1548109. 1574358. -3900. -8750. Actual: 1628641. 1548187. 1574504. -3899. -8726. SETMIR VALUES SecDesOrient: 1257.00 0.00 -25.00 0.00 146.27 SecOrient: 1256.99 0.05 -25.09 0.21 146.52 ---=== PROBLEMS IN DETAIL ===---